# Towards Reliable Human Evaluations in Gesture Generation: Insights from a Community-Driven State-of-the-Art Benchmark

### v2

@@ 3.4 Takeaways
Our analysis in Sec. 3 reveals three profound shortcomings of evaluation practices in gesture generation:
@@
    Doubtful ecological validity of evaluations of multimodal grounding due to the uncontrolled confounding factor of motion realism.
@@ 4 Evaluation Protocol for the BEAT2 Dataset
Motivated by the findings of Sec. 3, we develop a human evaluation protocol for automatic 3D hand- and body gesture-generation models on the widely-used BEAT2 dataset. Our goals are to facilitate standardised evaluations, and to provide a template that can be easily extended to new datasets and modelling problems. Later, in Sec. 5, we validate our methodology by conducting a community-driven benchmarking of six competing models.
4.1 Evaluation Segment Selection and Rendering
Gesture generation aims for realistic and expressive animation rather than replication of the dataset. This distinction is important due to pose estimation artifacts and the natural variation of human expression, and necessitates careful selection of evaluation segments. Overlooking this step, as most evaluations in our survey do, can lead to lower scores for the reference human motion, and ultimately, imprecise results.
We curate 108 evaluation segments from the BEAT2 English test set, covering all speakers. The segments are randomly sampled complete sentences (according to the text transcription), manually filtered for artifacts like flickering and self-intersection. We selected four segments for most speakers, and eight for Scott and Wayne due to their higher mocap quality. This larger segment count, compared to typical evaluations (e.g., GENEA Challenges) allows for more reliable user studies whilst leaving room for analysis on the stimulus level.
To create the video stimuli shown to the crowdsourced evaluators, we develop a Blender-based rendering pipeline.To avoid retargeting errors, we directly visualise the SMPL-X models [63], with added textures for increased realism and the face hidden by a mask (as the unrelated facial motion may distract evaluators). For more details, please refer to Sec. A.8.
4.2 Evaluation Tasks
Following general practice (Sec. 3), we adopt the two evaluation tasks of motion realism and speech-gesture alignment. We construct a separate user study for each of the following evaluation questions:
  1. 1.
    Motion Realism: “In which video does the character gesture more like a real person?”.
  2. 2.
    Speech-Gesture Alignment: “In which video do the character’s movements fit the speech better?”
In contrast to common practice, we mute the audio during motion-realism evaluations. This important design choice allows us to evaluate the visual quality of synthetic gestures in complete isolation from how well they’re aligned with the speech.
4.3 User-Study Methodology
In order to reduce the cognitive load of crowdsourced evaluators, we propose only using pairwise comparisons for evaluation. While potentially less time efficient than direct ratings (since each user response involves watching two video stimuli instead of just one), pairwise tests lead to higher inter-rater reliability [76], which is crucial given the inherent difficulty of evaluating co-speech gesturing.
We include five response options on Likert-type scale (weak- or strong preference in either direction, and a tie). To capture detailed user feedback in an economical manner, we adapt the JUICE methodology [32] – originally for video generation – to 3D motion evaluation for the first time. JUICE (“JUstify their choICE”) requires test-takers to not only indicate their preferences, but also select pre-defined reasons for their choices.
In the next two sections, we describe the disentangled methodologies for our two evaluations. We refer the reader to the Appendix for implementation details such as precise question formulations, response options, JUICE factors, participant recruitment, attention checks, and more.
@@ 4.5 Speech-Gesture Alignment Methodology
Our literature review identified inaccurate multimodal alignment evaluation as a critical shortcoming (Sec. 3.1). The fundamental problem is that direct multimodal alignment ratings can be strongly confounded by the visual appeal of the motion, failing to distentangle the two aspects. We advocate for the mismatching methodology, introduced by the GENEA Challenges, as an effective solution.
Previous mismatching studies presented video clips with identical speech but different motion – one matched, one mismatched with the audio. However, Kucherenko et al. [45] found that differences in human-likeness ratings between individual motion segments (regardless of whether they were matched to the speech) explained a significant fraction of user preferences also in mismatching studies. To address this, we propose audio mismatching, where each video in a pair has the exact same motion, but different speech audio: one matched, and one mismatched from another evaluation segment. This distinguishes our evaluation from prior works [38, 68, 87, 44, 45, 86].
By keeping motion constant between compared video clips, the realism of the motion can no longer bias the evaluators. Unlike all prior gesture-generation evaluations, this setup therefore fully disentangles the evaluation of motion realism (where only the visuals differ between stimulus pairs) from the evaluation of speech-gesture appropriateness (where the visuals are identical between the two stimuli in a pair).
We note that there remains a potential bias between different speech segments, e.g., due to different voices being perceived as more appealing than others. We eliminate this bias by always using the same speaker voice in both stimuli. We further remove any systematic effect of possible preferences for individual speech segments within a speaker by ensuring that every speech segment appears equally often in matched as in mismatched videos.
@@ 5 Experiments
To address the gap in the community’s understanding of the state of the art in gesture generation (Sec. 3), and to validate our human evaluation protocol (Sec. 4), we conduct a community-driven benchmarking effort. Due to known challenges with reproducibility in gesture generation, we invited authors of published models to participate by following our protocol, retraining their system under compatible settings, and submitting generated outputs for the full test set. In total, we compared six recently published gesture- generation models – DiffuseStyleGesture [80]; Semantic Gesticulator [89]; ConvoFusion [57]; RAG-Gesture [58]; AMUSE [21]; and HoloGest [20] – to each other and to the BEAT2 mocap recordings. For detailed information regarding the key features of these six systems, as well as a specification of how the retrained systems deviated from published models, please refer to Appendix B. We also experiment with JUICE factors and report those results in Sec. A.6 and Sec. A.7.
Additional results on automated metrics and their correlation with subjective human ratings are provided in Appendix C.
@@ 5.1 Motion-Realism Benchmark
We report the benchmark results obtained by following the motion realism protocol outlined in Sec. 4.4, and contrast them to prior results drawn from the status quo evaluation practices. We report Elo ratings with bootstrapped 95% confidence intervals; for statistical details, please refer to Sec. A.4.
@@ 5.2 Prior Results on Motion Realism
ConvoFusion and RAG-Gesture were evaluated with win rates in pairwise comparison, being preferred 43–47% percent of the time against human motion [57, 58]. HoloGest used a five-point Likert-type rating scale in an evaluation with 30 test takers and reported average human-likeness scores of 4.61 for natural motion, 4.47 for HoloGest, and 3.70 for DiffuseStyleGesture [20]. The user study in Semantic Gesticulator [89] did not find their system to be significantly different from human motion, whereas AMUSE [21] did. DiffuseStyleGesture received an average score of about 4.1 on a five-point scale, statistically indistinguishable from natural motion at 4.2 in their user study [80].
@@ 5.3 Our Results on Motion Realism
On Fig. 4, we report the Elo ratings acquired from our motion realism protocol (Sec. 4.4). The Mocap condition is at the top of the results with a mean Elo rating of 1133, setting an empirical upper bound for motion realism in this study. Amongst the machine-learning systems, ConvoFusion and RAG-Gesture rank as the strongest models with Elo ratings of 1102 and 1088, respectively. It makes sense for these systems to exhibit similar performance, given that RAG-Gesture builds on ConvoFusion. HoloGest follows closely behind RAG-Gesture with a mean Elo rating of 1084, slightly ahead of Semantic Gesticulator with an Elo rating of 1070. All four of these systems remain well within a performance band that is arguably consistent with the broader state of the art.
At the lower end of the scale, AMUSE and DiffuseStyleGesture display notably lower Elo rating at 824 and 701, respectively. Notably, while AMUSE and DiffuseStyleGesture exhibit much wider confidence intervals, this does not imply that the comparative performance of these system is less predictable. Rather, this is a consequence of early stopping in our adaptive evaluation due to their clear separation from other conditions.
Our results found a considerable gap between the best-performing gesture-generation models and AMUSE and DiffuseStyleGesture, even though both those systems originally reported state-of-the-art performance. This may be taken as evidence that that previous claims of high motion quality may not hold up under careful evaluation, or that faithfully adapting published models to a new dataset or to new dataset splits can be a significant challenge.
More importantly, comparing the top systems, we find that:
@@ 5.4 Speech-Gesture Alignment Benchmark
We also conduct the speech-gesture alignment evaluation as outlined in Sec. 4.5. We report bootstrapped 95% confidence intervals for audio mismatching scores, which are the weighted preference accuracy for the matched stimulus, independently computed for each condition, with strong preference votes counting twice. For more details, please refer to Sec. A.5.
Figure 5: Results of the speech-gesture appropriateness user study. Ratio of user preference towards matched over mismatched stimuli for each condition considered, and 95% confidence intervals acquired from bootstrapping. Conditions are ordered by estimated appropriateness score in descending order.
@@ 5.4.2 Our Results on Speech-Gesture Alignment
As shown in Fig. 5, our speech-gesture alignment study exposes significant differences compared to prior published evaluations.
As expected, Mocap set the empirical topline with $\approx 74\%$ mean audio mismatching score. This far outperforms all generative models. DiffuseStyleGesture and HoloGest performed the best amongst the synthetic conditions, both scoring $\approx 60\%$. The near-identical alignment results suggest that HoloGest’s previously reported alignment advances over DiffuseStyleGesture may in fact be confounded by motion quality, as postulated in Sec. 3.1.
Semantic Gesticulator reached $\approx 57\%$, sharply contrasting prior results indicating strong semantic alignment. AMUSE, ConvoFusion, and RAG-Gesture all have confidence intervals overlapping $50\%$, indicating that their speech-gesture alignment of is not statistically significantly different from chance rate without FDR correction. This can be taken as evidence that their output could be interchanged for random gesture sequences in this task (or at least for random gesture sequences aligned to the start of a sentence, since that is how our segments were selected).
@@
The use of mismatching, as featured here, is crucial in disentangling motion quality from appropriateness for speech. As it stands, the results in current literature are not reliable and convey a false sense of progress, where in reality we have very far left to go.
@@ 6 Limitations
Our conclusions for state-of-the-art performance – i.e., that motion realism is nearing saturation, while speech-gesture alignment is far from a solved problem – is naturally limited by the scale and quality of BEAT2, and may not extend to other datasets. Adapting our evaluation protocol to other datasets is an important future direction.
The use of Elo ratings [28] when evaluating motion realism comes with an assumption that the collective preferences of the test takers are transitive, which need not be the case; cf. Arrow [7]. Boubdir et al. [11] provides a discussion of this assumption in Elo-ratings-based evaluation of machine-learning models, along with possible alternatives. That said, the vast majority of existing gesture evaluations surveyed in Sec. 3 also assume transitivity, e.g., by having test takers assign ordinal ratings to different systems.
For speech-gesture alignment, we want to mention that the mismatching-based evaluation really measures specificity more than inherent appropriateness. If a gesture-generation system were to generate hypothetical “universally appropriate outputs” $x_{\mathrm{universal}}$ that are a good fit to any possible speech audio segment $s$, such a system will score close to chance level in a mismatching-based appropriateness study. In practice, human motion contains rhythms at multiple levels [66], so a given motion $x$ can be perceived as rhythmically appropriate to a range of different speech audio with different rhythms (cf. Miller et al. [56], Avirgan [8]), whereas semantic gestures (if produced by a system) are likely to be much more specific to a given context and speech. This may, among other things, cause our evaluations to be relatively more sensitive to semantic gestures and their appropriateness over rhythmic (beat) appropriateness. On the other hand, beats are vastly more common than semantic gestures, and are overall still likely to influence appropriateness evaluations more; cf. Saund and Marsella [70].
@@
We also note that some of the systems in our evaluation required adaptations, and re-training, for the BEAT2 dataset. Although we mitigated these effects by having the original authors submit their outputs, such changes may nevertheless have an adverse effect on model performance, We hope that the authors of future models will perform first-party evaluation using our protocol and our released data, for more accurate measurement of the state-of-the-art.
Finally, the lack of reliable automated metrics is a significant problem. The human preference responses collected during our evaluation are suitable for training human opinion predictors [35], and may also enable the validation of existing and new automated metrics [25, 37].
@@ 7 Conclusion
In this paper, we performed a detailed survey of the various systematic problems in gesture-generation evaluations, declaring that it may not be possible to determine the state of the art from published results and that the lack of standardisation may contribute to a false sense of progress. We then designed a new evaluation protocol, based on best practices from prior evaluations, with added innovations (Elo ratings, audio mismatching, JUICE) for improved comparability and nuance without confounding. Finally, we put our protocol to the test by benchmarking six recent gesture-generation models, all of which have reported state-of-the-art performance.
Our evaluation results provide a first-of-its-kind benchmark between competing already published models, in a field severely lacking direct comparisons. We found strong evidence that motion realism on BEAT2 is no longer a distinctive factor between state-of-the-art models. Perhaps even more importantly, we show how current evaluation practices may lead to over-inflated multimodal-alignment results due to entanglement between evaluation dimensions. In contrast, our disentangled evaluations avoid common pitfalls, and pave the way towards more specialised evaluations for, e.g., semantic alignment or emotional expression.
Finally, we will release all collected data from our evaluations, including 16000 human votes, 750 rendered video segments, and five hours of synthetic motion from the benchmarked models. We hope that the proposed evaluation protocol and the released data will pave the way towards standardised gesture-generation evaluations [59] on currently used datasets, and enable the development of more targeted human evaluation methodologies and better automated metrics.
8 Acknowledgements
@@
Appendix A User study set up and analysis
We now describe in more detail the experience of taking one of the user study, and how test takers were recruited.
A.1 Study participants

### v4

@@ 3.4 Takeaways
Overall, we identified three critical shortcomings in current human evaluation practices in gesture generation:
@@
    Doubtful ecological validity of evaluations of multimodal alignment, due to the uncontrolled confounding factor of motion realism.
@@ 4 Evaluation Protocol for the BEAT2 Dataset
Motivated by the findings of Sec. 3, we develop a human evaluation protocol for automatic 3D hand- and body gesture-generation models on the widely-used BEAT2 dataset [47]. Our goals are to establish standardised evaluations, and to provide a template that can be easily adapted and extended for new datasets and modelling problems.
Using the GENEA Challenges as a starting point, our proposed protocol offers several methodological improvements to human evaluation of gesture generation. Later, in Sec. 5, we use the proposed protocol to conduct a community-driven benchmark of six recent models.
4.1 Disentangled Evaluation Dimensions
Following general practice (cf. Sec. 3), we choose motion realism and speech-gesture alignment as the two evaluation dimensions of our protocol. Our review found that current human evaluation practices may lead to highly correlated measurements between the two dimensions (cf. Sec. 3.1). This means that prior evaluations of methods claiming to improve, e.g., semantic gesturing, or temporal alignment, are unreliable due to the significant confounding effect of motion realism. To combat this problem, we propose a disentangled methodology similar to the GENEA Challenges [42, 43, 44].
First, during motion-realism evaluations, we mute the audio in order to assess the visual quality of synthetic gestures in complete isolation from how well they’re aligned with the speech. Second, to measure speech-gesture alignment, we employ a mismatching methodology where every model is only compared against its own (mismatched) outputs, completely removing the confounding effect of motion realism. Overall, these important design choices ensure that the two evaluation dimensions are fully independent.
We remark that motion realism and speech-gesture alignment are not truly separable in human perception of gesturing. Whether motion looks natural depends on the context, and motion quality must be at a high enough level for motion to be appropriate. Nevertheless, it is crucial to separate the two aspects when evaluating synthetic gestures, because they are often targeted by distinct modelling choices.
@@ 4.2 Motion-Realism Methodology
To evaluate motion realism, we pose the following question to user-study participants: “In which video does the character gesture more like a real person?”.
In order to reduce the cognitive load of crowdsourced evaluators, we propose only using pairwise comparisons as they lead to higher inter-rater reliability [74], which is crucial given the inherent difficulty of evaluating co-speech gesturing. For each pairwise comparison, we include five response options on a Likert-type scale (weak- or strong preference in either direction, and a tie). However, to capture detailed user feedback in an economical manner, we adapt the JUICE methodology [31] – originally for video generation – to 3D motion evaluation for the first time. JUICE (“JUstify their choICE”) requires test-takers to not only indicate their preferences, but also select pre-defined reasons for their choices.
4.2.1 Rating System
@@ 4.3 Speech-Gesture Alignment Methodology
To evaluate speech-gesture alignment, we develop a mismatching protocol with the user-study question: “In which video do the character’s movements fit the speech better?”.
Similar to the motion-realism study, we combine five response options with additional JUICE feedback options detailed in the Appendix. Previous mismatching studies [44, 43] presented video clips with identical speech but different motion – one matched, one mismatched with the audio. However, Kucherenko et al. [44] found that differences in motion-realism ratings between individual motion segments (regardless of whether they were matched to the speech) explained a significant fraction of user preferences also in mismatching studies. To address this, we propose audio mismatching, where each video in a pair has the exact same motion, but different speech audio: one matched, and one taken from another evaluation segment. This distinguishes our evaluation from prior works [37, 66, 85, 43, 44, 83].
By keeping motion constant between compared video clips, the realism of the motion can no longer bias the evaluators. Unlike all prior gesture-generation evaluations, this setup therefore fully disentangles the evaluation of motion realism (where only the visuals differ between stimulus pairs) from the evaluation of speech-gesture appropriateness (where the visuals are identical between the two stimuli in a pair). We note that there remains a potential bias between different speech segments, e.g., due to different voices being perceived as more appealing than others. We eliminate this bias by always using the same speaker voice in both stimuli. We further remove any systematic effect of possible preferences for individual speech segments within a speaker by ensuring that every speech segment appears equally often in matched as in mismatched videos.
4.4 Additional details
Please refer to Appendix B for more information on our protocol, including test segment selection, user-study screens, crowdsourcing design, 3D visualisation, and the proposed JUICE factors for each evaluation.
@@ 5 Experiments
To address the gap in the community’s understanding of the state of the art in gesture generation (Sec. 3), and to validate our human evaluation protocol (Sec. 4), we conduct a community-driven benchmarking effort. Due to known challenges with reproducibility in gesture generation, we invited authors of published models to participate by following our protocol, retraining their system under compatible settings, and submitting generated outputs for the BEAT2 English test set. To prevent cherry picking, we collect five sets of samples from each model (corresponding to five random seeds) for each clip in the test set.
In total, we compared six recently published gesture- generation models – DiffuseStyleGesture [78]; Semantic Gesticulator [87]; ConvoFusion [57]; RAG-Gesture [56]; AMUSE [19]; and HoloGest [18] – to each other and to the BEAT2 mocap recordings, recruiting over 600 evaluators and collecting over 16,000 pairwise votes on the Prolific crowdsourcing platform. For detailed information regarding the evaluated systems, test-taker demographics, user-study construction, JUICE analysis, and automatic metrics, we refer the reader to the Appendix.
@@ 5.1 Motion-Realism Benchmark
We report the benchmark results obtained by following the motion realism protocol outlined in Sec. 4.2, and contrast them to prior results drawn from the status quo evaluation practices.
@@ 5.2 Prior Results on Motion Realism
ConvoFusion and RAG-Gesture were evaluated with win rates in pairwise comparison, being preferred 43–47% percent of the time against human motion [57, 56]. HoloGest used a five-point Likert-type rating scale in an evaluation with 30 test takers and reported average human-likeness scores of 4.61 for natural motion, 4.47 for HoloGest, and 3.70 for DiffuseStyleGesture [18]. The user study in Semantic Gesticulator [87] did not find their system to be significantly different from human motion. AMUSE [19] outputs were tied with or preferred to motion capture 49% of the time. DiffuseStyleGesture received an average score of about 4.1 on a five-point scale, statistically indistinguishable from natural motion at 4.2 in their user study [78].
@@ 5.3 Our Results on Motion Realism
On Fig. 4, we report Elo ratings with bootstrapped 95% confidence intervals acquired from our motion realism protocol (Sec. 4.2). The Mocap condition is at the top of the results with a mean Elo rating of 1133, setting an empirical upper bound for motion realism in this study. Amongst the machine-learning systems, ConvoFusion and RAG-Gesture rank as the strongest models with Elo ratings of 1102 and 1088, respectively. It makes sense for these systems to exhibit similar performance, given that RAG-Gesture builds on ConvoFusion. HoloGest follows closely behind RAG-Gesture with a mean Elo rating of 1084, slightly ahead of Semantic Gesticulator with an Elo rating of 1070. All four of these systems remain well within a performance band that is arguably consistent with the broader state of the art.
At the lower end of the scale, AMUSE and DiffuseStyleGesture display notably lower Elo rating at 824 and 701, respectively. While AMUSE and DiffuseStyleGesture exhibit much wider confidence intervals, this does not imply that the comparative performance of these systems is less predictable. Rather, this is a consequence of early stopping in our adaptive evaluation due to their clear separation from other conditions.
Our results found a considerable gap between the best-performing gesture-generation models and AMUSE and DiffuseStyleGesture, even though both of those systems originally reported state-of-the-art performance. This may be taken as evidence that previous claims of high motion quality may not hold up under careful evaluation, or that faithfully adapting published models to a new dataset or to new dataset splits can be a significant challenge. Comparing the top systems, we find that:
@@ 5.4 Speech-Gesture Alignment Benchmark
We also conduct the speech-gesture alignment evaluation as outlined in Sec. 4.3. We report bootstrapped 95% confidence intervals for audio mismatching scores, which are the weighted accuracies of preferring the matched stimulus, independently computed for each condition, with strong preference votes counting twice. For more details, please refer to the Appendix.
Figure 5: Results of the speech-gesture alignment user study. Ratio of user preference towards matched over mismatched stimuli for each condition considered, and 95% confidence intervals acquired from bootstrapping. Conditions are ordered by estimated appropriateness score in descending order.
@@ 5.4.2 Our Results on Speech-Gesture Alignment
As shown in Fig. 5, our speech-gesture alignment study exposes significant differences compared to prior published evaluations. As expected, Mocap set the empirical topline with ${\approx}74\%$ mean audio mismatching score. DiffuseStyleGesture and HoloGest performed the best amongst generative models, both scoring ${\approx}60\%$. The near-identical alignment results suggest that HoloGest’s previously reported alignment advances over DiffuseStyleGesture may in fact be confounded by motion quality, as postulated in Sec. 3.1. Semantic Gesticulator reached ${\approx}57\%$, sharply contrasting prior results indicating strong semantic alignment. AMUSE, ConvoFusion, and RAG-Gesture all scored near 50% mismatching scores, meaning that the motion they generate is only as valid for their input speech as for randomly chosen speech segments.
@@
The use of mismatching, as featured here, is critical in disentangling the quality of motion from its alignment with the speech. Prior evaluations of speech-gesture alignment are not reliable, and convey a false sense of progress, where in reality we have very far left to go.
@@ 6 Limitations
Our conclusions for state-of-the-art performance – i.e., that motion realism is nearing saturation, while speech-gesture alignment is far from a solved problem – are naturally limited by the scale and quality of BEAT2, and may not extend to other datasets. Adapting our evaluation protocol to other datasets is an important future direction.
Our mismatching-based evaluation of speech-gesture alignment is a diagnostic measure of specificity. If a gesture-generation system were to generate hypothetical “universally appropriate outputs” $x_{\mathrm{universal}}$ that are a good fit to any possible speech audio segment $s$, such a system will score close to chance level in a mismatching-based appropriateness study. In practice, human motion contains rhythms at multiple levels [64], so a given motion $x$ can be perceived as rhythmically appropriate to a range of different speech audio with different rhythms (cf. [55, 8]), whereas semantic gestures (if produced by a system) are likely to be much more specific to a given context and speech. This may, among other things, cause our evaluations to be relatively more sensitive to semantic gestures and their appropriateness over rhythmic (beat) appropriateness. On the other hand, beats are vastly more common than semantic gestures, and are overall still likely to influence appropriateness evaluations more; cf. Saund and Marsella [68].
@@
We also note that some of the systems in our evaluation required adaptations, and re-training, for the BEAT2 dataset. Although we mitigated these effects by having original authors prepare the submissions, such changes may nevertheless have an adverse effect on model performance, We hope that the authors of future models will perform first-party evaluation using our protocol and our released data, for more accurate measurement of the state of the art.
Finally, the lack of reliable automatic metrics is a significant problem. The human preference responses collected during our evaluation are suitable for training human opinion predictors [34], and may also enable the validation of existing and new automatic metrics [23, 36].
@@ 7 Conclusion
In this paper, we surveyed systematic problems in gesture-generation evaluations and found that the state of the art cannot be determined from published results, and that the lack of standardisation may contribute to a false sense of progress. We then designed a new evaluation protocol, innovating on prior best practices with Elo ratings, audio mismatching, and JUICE. Finally, we put our protocol to the test by benchmarking six recent gesture-generation models, all of which have reported state-of-the-art performance.
Our results provide a first-of-its-kind benchmark between competing published models in a field severely lacking direct comparisons. We find that motion realism on BEAT2 is no longer a distinctive factor between leading models, and – perhaps even more importantly – we show that current evaluation practices inflate multimodal-alignment results due to entanglement between evaluation dimensions. In contrast, our disentangled evaluations avoid common pitfalls, and pave the way towards reliable evaluations of more specialised aspects, e.g., semantic alignment or emotional expression.
We release all data collected in our evaluations (16,000 human votes, over 750 rendered video segments, and five hours of synthetic motion from the benchmarked models) to support standardised gesture-generation evaluations and the development of more targeted human evaluation methodologies and better automatic metrics.
Acknowledgements
@@
Appendix A Additional Results
Since this paper was submitted, we have extended our benchmark with an additional model from the authors of the Seamless Interaction dataset [2]. The Seamless model is a 250M-parameter Diffusion Transformer [62] trained with a conditional flow-matching objective on the Seamless Interaction dataset, which we evaluated in a zero-shot generalisation setting using the BEAT2 test set under the same conditions as before. As part of this extended evaluation, we also collected new votes for previous systems and system pairs, altogether extending our public database of human preference ratings from 16,000 to 20,000 votes.
We present the extended evaluation results in Fig. 6 and Fig. 7. To our surprise, the Seamless model not only matched BEAT2 motion capture in terms of motion realism, but also achieved near-indistinguishable appropriateness scores in our audio-mismatching evaluation. This is a breakthrough result that underlines the importance of evolving datasets and standardised evaluations for tracking the state-of-the-art. We emphasise, however, that these high ratings do not imply that the generated motion is perfect: the BEAT2 dataset itself contains motion-capture artifacts, and its actors exhibit varying levels of expressivity and alignment in their gestures. Similarly, our mismatching evaluation is just a first (albeit important) step towards measuring multimodal alignment: there is a strong need for reliable evaluation methods for measuring, e.g., semantic or emotional alignment of gestures.
Figure 6: Updated Elo ratings after adding the Seamless model to the benchmark.
Figure 7: Updated appropriateness scores after adding the Seamless model to the benchmark.
Appendix B Additional Details on our Evaluation Protocol
B.1 Test Segment Selection
Following prior large-scale evaluations in speech [63] and gestures [42, 44, 43], we use short speech-gesture segments as the stimuli for our user studies. Using this one set of speech segments as the basis for all user studies carried out on the BEAT2 dataset, we control for the effect of the speech on system behaviour (e.g., the same speakers are always represented in the same way in every evaluation).
In particular, we curate 108 evaluation segments from the BEAT2 English test set, covering all speakers. The segments are randomly sampled complete sentences, manually filtered for artifacts like flickering and self-intersection. We selected four segments for most speakers, and eight for Scott and Wayne due to their higher mocap quality. This larger segment count, compared to typical evaluations (e.g., GENEA Challenges) allows for more reliable user studies whilst leaving room for analysis on the stimulus level. The criteria for selecting speech segments for the user studies were as follows:
  • Each segment should correspond to one or more complete sentences.
  • Segment duration should be within the range of 7.0 to 12.0 seconds.
  • Segments should be disjunct (no overlap).
Although a lot of the test dataset contains somewhat awkward finger poses due to the difficulty of tracking fingers, this was deemed less visually distracting and would be more drastic to exclude, so it was not considered grounds for exclusion.
The only cases where mesh penetration were permitted were (2a) when the penetration visually resembled clothing or tissue giving way to light pressure, or (2b) where the penetration occurred due to the aforementioned poor finger tracking, as long as the fingers at worst were merely seen clipping into each other, and not passing through each other to the other side at any point. We decided to retain segments satisfying (2b) for the evaluations because these instances of mesh penetration are associated with poses having gestural importance, such as interwoven hands or finger against palm, which are important not to exclude.
After the segment selection was performed, we measured the potential rhythm bias of our alignment evaluation by comparing our evaluation segments to the semantic labels of BEAT2; we found that ${\approx}59\%$ of segments fully contain at least one semantic gesture, even though we did not intentionally select segments with semantic annotations. In other words, our benchmark heavily oversamples semantic gestures compared to their natural frequency, and is not unnecessarily unfair towards models aiming at semantic generation.
Overall, gesture generation aims for realistic and expressive animation rather than replication of the dataset. This distinction is important due to pose estimation artifacts and the natural variation of human expression, and necessitates careful selection of evaluation segments. Overlooking this step, as most evaluations in our survey do, can lead to lower scores for the reference human motion, and ultimately, imprecise results.
B.2 3D Visualisation
Following the approach of the GENEA Challenges [85, 44], our visualisation (Fig. 8) includes full-body motion with root-node translation and rotation. This captures the character’s positioning and stance with respect to the camera, leading to a more lifelike and expressive visualisation than methods that restrict animation to only the upper body [42, 87, 60, 19].
Since BEAT2 is in the SMPL-X format, we use SMPL-X meshes for our visualisation, which offer an anatomically accurate body shape and proportion that matches each speaker [61]. Although SMPL-X includes gendered meshes, we opted to use the gender-neutral mesh for all characters, which is further modified by the per-speaker body shape parameters. This decision was made because the gendered meshes often resulted in increased self-intersection when visualising motion from the BEAT2 test set. However, the gender-neutral mesh lacks features such as hair and clothing and does not replicate realistic skin tones or facial characteristics. The male- and female-presenting speakers are distinguished through different SMPL-X textures: male characters use the default white shirt texture, and female characters use the pink one, matched to the perceived gender of the voice. These textures also introduce variation in skin tone. To enhance realism, we added a hair prop and applied a displacement map to the clothing, improving its appearance and reducing the flatness seen in the SMPL-X mesh. Due to the absence of gaze information and inaccurate lip sync in BEAT2, we covered the face with a white mask. This avoids distracting visual artefacts that could negatively influence test-taker perception.
Furthermore, we excluded the feet from view, cutting the frame at approximately knee level. This was a deliberate choice motivated by known issues with foot-ground interactions, such as sliding and ground penetration, which are both common in synthetic motion and easily detectable by human observers. If shown, these artefacts would likely overshadow the gestural qualities under evaluation, as crowdsourced participants tend to focus on the most visually salient errors (see, for instance, the results in Sec. C.3). By omitting the feet, we help raters concentrate on the gestures themselves, aligning with the primary evaluation objective.
It is also worth noting that while foot-ground contact issues are primarily governed by straightforward laws of physics rules and can potentially be resolved through post-processing, upper-body gesturing is much less dictated by physical laws and is instead rooted in communicative and cultural conventions. This arguably makes gesturing a deeper and more challenging problem to solve in the long term, compared to issues of character interaction with the ground plane. For the record, while the human evaluations are conducted based on cropped visualisations, the automatic metrics described in Appendix E operate on the full-body pose and motion data.
B.3 User-Study Setup
For the user-study screens, we propose the layout and phrasing (for instructions and response options) shown in Fig. 9. Our implementation of this interface, used for the experiments in Sec. 5, is illustrated in Fig. 10. We suggest presenting each crowdsourced participant with 25 screens, leading to 25 pairwise votes collected in total (barring technical issues).
B.3.1 Details on Participant Recruitment
As participant recruitment may highly depend between evaluation setups, we do not aim to standardise it in our protocol. Regardless, we share important details from our evaluations in Sec. 5 below.